OpenStudy (anonymous):

Solve the complex variable equation (-9/2)*sqrt(3) + (9/2)*i So that it is in the form ______*exp(i_______)

OpenStudy (anonymous):

you need two numbers \(r\) and \(\theta\) \(r=\sqrt{a^2+b^2}\) which in your case is \[\sqrt{\left(-\frac{9\sqrt3}{2}\right)^2+\left(\frac{9}{2}\right)^2}\]

OpenStudy (anonymous):

then \(\theta\) find via \(\sin(\theta)=\frac{b}{r}\) and \(\cos(\theta)=\frac{a}{r}\) sometimes you can use \(\theta =\tan^{-1}(\frac{b}{a})\) but not in this case because you are in the second quadrant